Circular No. 6513

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

EG CANCRI P. Schmeer, Bischmisheim, Germany, reports that this probable SU

UMa-type dwarf nova is in outburst for the first time since 1977 Nov., when it reached m_{pv} = 11.9 (IBVS 2401, 2926). E. Waagen, American Association of Variable Star Observers, reports that AAVSO data confirm no outbursts since the early 1980s, when the star was added to their program. The following visual magnitude estimates are partly from Waagen: Nov. 21.082 UT, [15.2 (G. Poyner, Birmingham, England); 21.23, [13.7 (P. Schmeer, Bischmisheim, Germany); 30.917, 12: (Schmeer); 30.951, 12.0 (Schmeer); 30.989, 11.8 (Schmeer); Dec. 1.021, 12.1 (T. Vanmunster, Landen, Belgium); 1.248, 11.7 (Schmeer); 2.298, 12.0 (G. Hanson, Cave Creek, AZ); 2.333, 11.4: (T. Burrows, Novato, CA).

SUPERNOVA 1996bv IN UGC 3432 M. Cavagna, Sormano, Italy, report the following precise position of

SN 1996bv, obtained with the 0.5-m reflector (+ CCD) on Nov. 22.968 UT: α = 6^{h}16^{m}13^{s}.00, δ = +57^{o}03′08′′.9 (equinox 2000.0). D. Balam, University of Victoria, reports a preliminary magnitude of V = 15.5 ± 0.4, obtained on Nov. 23 with the 1.82-m Plaskett telescope at the Dominion Astrophysical Observatory (National Research Council of Canada), and the following position end figures: 13^{s}.04, 08′′.4.

COMET C/1995 O1 (HALE-BOPP) Total (visual) magnitude and coma-diameter estimates: Nov. 5.73 UT,

4.9, – (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars); 8.75, 4.7, ∼ 15′ (M. V. Zanotta, San Marco Pass, Italy, naked eye); 10.06, 4.6, – (A. Hale, Cloudcroft, NM, naked eye); 20.06, 4.6, – (R. Keen, Mt. Thorodin, CO, naked eye); 27.72, 4.5, 12′ (B. H. Granslo, Tryvann, Oslo, Norway, 10×50 binoculars); 29.78, 4.4, 9′ (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); Dec. 1.71, 4.2, ∼ 35′ (H. Mikuž, Ljubljana, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + CCD).

COMET C/1996 Q1 (TABUR) Total visual magnitude and coma-diameter estimates by K. Hornoch,

Lelekovice, Czech Republic (B = 10×80 binoculars; L = 0.35-m reflector): Nov. 6.72 UT, 8.7, 7′ (B); 8.73, 9.6:, 12′ × 6′ (B); 23.18, 9.4, 7′ (L).

1996 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green